OpenStudy (anonymous):

Derivative of y= x/square root of 1+x^2

OpenStudy (anonymous):

This should remind you of the pythagorean theorem.

OpenStudy (anonymous):

Try drawing the appropriate right triangle.

OpenStudy (anonymous):

This has nothing to do with pythagorean theorem, it's calculus (Derivatives)

OpenStudy (anonymous):

Oh, then you can completely ignore this then:  I'm sure \(\large \frac{x}{\sqrt{1+x^2}}\) couldn't possibly have anything to do with the sine function.

OpenStudy (anonymous):

using quotient rule y' = \[\sqrt{1+x^2} - 2x^2(\sqrt{1+x^2} / 1+x^2\]

OpenStudy (anonymous):

the whole thing is divided by 1+x^2 and the quotient rule is this [f'(x)g(x) - f(x)g'(x)] / g(x)^2

OpenStudy (anonymous):

Or . . . y = sin(Θ) y' = cos(Θ)

OpenStudy (anonymous):

 See what I'm getting at here @Mcisn3 , @jayz657 ?

OpenStudy (anonymous):

quotient rule for this is easy enough so that you do not need a trig sub to do this

OpenStudy (anonymous):

Thanks jayz657, that makes sence :)